OpenStudy (anonymous):

solve and show me the working out of 3(x-2) = -2(x+19)

OpenStudy (anonymous):

x = -32/5

OpenStudy (anonymous):

3(x-2)=-2(x+19) 3x-6=-2x-38 5x=-32 x=-32/5

OpenStudy (anonymous):

distribute the 3 on the left to the parenthesis: 3x - 6 = -2 (x + 19) Now distribute the 2 on the right side, to the parenthesis 3x - 6 = -2x - 38 Then add -2x to the other side to combine like terms. 3x - 6 = -2x - 38 +2x +2x 5x -6 = -38 Then add 6 to -38, in order to add like terms. 5x -6 = -38 +6 +6 5x = -32 Then divide 5 from each side to solve for x 5x = -32 /5 /5 Then you get: x = -32/5